OpenStudy (anonymous):

Find dx/dy of y=(x+sqrt(x))^-2

OpenStudy (imstuck):

Are you finding dx/dy or dy/dx?

OpenStudy (anonymous):

dy/dx sorry, miss type

OpenStudy (freckles):

Did you apply chain rule?

OpenStudy (anonymous):

I don't understand how to use the chain rule

OpenStudy (freckles):

what if you were to just differentiate x^(-2) w.r.t x?

OpenStudy (freckles):

could you do that

OpenStudy (anonymous):

-2x^-3

OpenStudy (freckles):

\[\frac{dy}{dx}=-2(x+\sqrt{x})^{-3} \cdot (x+\sqrt{x})'\] So now you just need to find the derivative of the inside

OpenStudy (freckles):

\[\frac{d}{dx}(x+\sqrt{x})=?\]

OpenStudy (freckles):

\[\frac{d}{dx}(x+\sqrt{x})=\frac{d}{dx}(x+x^\frac{1}{2})=?\]

OpenStudy (anonymous):

would it be...\[1+\frac{ 1 }{ 2\sqrt{x} }\]

OpenStudy (freckles):

yes

OpenStudy (freckles):

\[\frac{dy}{dx}=-2(x+\sqrt{x})^{-3} \cdot (x+\sqrt{x})' \\ \frac{dy}{dx}=-2(x+\sqrt{x})^{-3} (1+\frac{1}{2 \sqrt{x}}) \]

OpenStudy (anonymous):

ok thank you!